CLASSICAL, SEMI-CLASSICAL, AND QUANTUM DYNAMICS OF UNI-AXIAL AND MULTI-AXIAL FLOPPY ROTORS

\underline{WILLIAM G. HARTER} and HORACE T. CROGMAN, Department of Physics, University of Arkansas, Fayetteville, AR 72701.

Molecules carrying more or less freely turning rotors have classical and quantum rotational properties that are more difficult to calculate and visualize than those of a more conventional semi-rigid molecule.^{a} The analogy with a gyro-stabilized spacecraft provides a starting point for discussion, but the quantum mechanics of angular momentum provides a far richer state space than the corresponding classical model. To help unravel the many dynamical and spectral possibilities, a semi-classical analysis involving rotational energy surfaces RES^{b} may be used to elucidate both classical and quantum modeling. We consider the simplest models of a single uniaxial rotor, such as a methyl group, attached to a larger rigid or semi-rigid molecule, and compare RES geometry to quantum eigenvalue solutions. The possibility emerges for more easily analyzing multi-axial species that are connected in series or parallel to a central rotor.

aJ. Ortigoso and J. T. Hougen, J. Chem. Phys. 101, 15 (1994).

bW. G. Harter, Princ. of Symmetry, Dynamics & Spectroscopy (Wiley 1993) p. 608.